# Reward-Robust RLHF in LLMs

### v1

@@ 5.1 BRME: Uncertainty Set Modeling
Figure 2: Diagram for BRME. Each head outputs the mean and the standard deviation of the corresponding reward distribution and reparametrization is emplyed to address the non-differentialbility.
In this section, we will show how we model the uncertainty set $\mathcal{R}^{\text{uncertain}}$ as well as the nominal reward function $\widehat{r}$. We propose Bayesian Reward Model Ensembles (BRME): We train a multi-head Bayesian reward model where the reward is modeled as a Gaussian distribution. Each head $i$ has two outputs: one representing the mean and the other representing the standard deviation. A sample from this distribution is then output as the reward.
The diagram for BRME is shown in Figure 2. To facilitate deployment and conserve computing resources, we do not train multiple reward models independently as an ensemble; instead, we use parameter sharing. All reward heads share a common base model, which serves as a feature extractor. During training, each data sample is randomly assigned to a head for training, with reparameterization employed to address the non-differentiability of the sampling process.
The training process is divided into two stages. In the first stage, we train a normal one-head reward model following the loss function in (3). In the second stage, we leverage a MSE loss (8) to train the reward model, which is first introduced in Wu et al. (2024). The use of MSE loss to train the ensembled reward model ensures that: 1) the output’s standard deviation reflects the confidence of the model, 2) the scoring range of each reward head is consistent. We defer the training pipeline to Appendix B.
During prediction, each sample is evaluated by all heads, and the reward output by the head with the smallest standard deviation is selected as the nominal reward. We provide the theoretical analysis to prove that the head with the smallest standard deviation provides the most confident result in Appendix B.
@@ 5.2.1 Experimental setup
We use LLaMa3-8B-Instruct Dubey et al. (2024) as the initialization of the actor. In BRME setting, we train a single-head RM and a 5-head BRME starting from LLaMa3-8B-Instruct Dubey et al. (2024) as is described in Appendix B. The single-head reward model is used as the initialization of the critic model and the ensembled reward model is used as the reward signal source.
@@
For the training process, we use HH-RLHF Bai et al. (2022a) and UltraFeedBack Cui et al. (2024) to train the BRME. HH-RLHF, UltraFeedBack Cui et al. (2024) , along with an internal prompt dataset collected by the PM team, is employed to implement the PPO algorithm. Details of the datasets are deferred to Table 6. For performance evaluation, we select ARC Clark et al. (2018), LAMBADA Paperno et al. (2016), PIQA Bisk et al. (2020), SciQ Welbl et al. (2017), WinoGrande Sakaguchi et al. (2019), TQA Lin et al. (2022), MMLU Hendrycks et al. (2020), GSM8K Cobbe et al. (2021), FDA Arora et al. (2023), EQ-Bench Paech (2023), Arithmetic Brown et al. (2020), and ANLI Nie et al. (2020) as our benchmarks. The evaluation dimensions include robustness, general knowledge, numerical computation, emotion reading, information extraction, reasoning, context understanding, and commonsense. Detailed descriptions of the datasets are shown in Table 7.
@@
The reward model training and PPO experiments are conducted on 24 Nvidia H800-SXM-80GB GPUs in 3 nodes using DeepSpeed library and Zero stage 2 Rasley et al. (2020), and HuggingFace Accelerate Gugger et al. (2022). In PPO process, the actor model and the critic model occupy 10 gpus respectively. The initialized SFT model and the reward model occupy 2 gpus respectively. We use AdamW optimizer Loshchilov et al. (2017). The experience batch size in PPO is set to be 128.
@@ 5.2.2 Main Results
We incrementally increased the performance trade-off hyperparameter $\lambda$ from 0 to 1 in intervals of 0.2 and repeated PPO training under each setting for 800 steps. The performance evaluation results at 200 and 800 steps are shown in Figure 3. It is important to note that when $\lambda=1$, the algorithm reduces to standard RLHF with a single nominal reward model.
@@
In the short run (200 steps), although there are a few exceptions—such as ANLI-r1 and LAMBADA, where $\lambda$ outperforms the other settings—in most cases, the trade-off versions with $\lambda=0.4$ and $\lambda=0.6$ show better performance. This suggests that, even early in training, incorporating a balance between performance and robustness offers notable advantages. The ability of the reward-robust RLHF framework to temper the optimization process appears to result in more stable performance gains compared to standard RLHF. Compared with standard RLHF, the accuracy of reward-robust RLHF increases by 0.99% and 1.40% respectively when $\lambda=0.4$ and $0.6$.
@@
Over the long run (800 steps), the advantages of incorporating robustness into the training process become even more pronounced. As shown in the results, nearly all experimental groups with $\lambda\neq 1$ exhibit improved performance by the end of 800 steps, confirming the long-term benefits of reward-robust RLHF. In contrast, standard RLHF with $\lambda=1$ not only fails to capitalize on the full optimization process but, in some cases, even results in negative performance growth. For example, in tasks like ANLI-r1 and GSM8K, we observe a decrease in model capability during the additional 600-step optimization process under standard RLHF. This highlights a key limitation of the purely performance-driven RLHF approach: its vulnerability to overfitting or misguidance from imperfect reward signals, leading to capability degradation over time. Compared with standard RLHF, the accuracy of reward-robust RLHF increases by 2.42% and 2.03% respectively when $\lambda=0.4$ and $0.6$.
@@
On tasks where the reward model performs poorly, such as MMLU, we notice that PPO training under standard RLHF can degrade model performance, evidenced by a drop in accuracy, which is reflected in the bar graphs pointing downward. This degradation likely results from the model being optimized based on unreliable or misleading reward signals, which causes the policy to drift away from optimal behavior. In contrast, the reward-robust RLHF approach, which balances performance with robustness ($\lambda=0.4$ or $\lambda=0.6$), mitigates this decline. By narrowing the optimization focus and stabilizing the reward signal, the model is better able to resist overfitting to faulty rewards, ultimately preserving and improving performance.
The results suggest that in settings where the reward model is imperfect or data is sparse, adopting a reward-robust approach can significantly improve training stability and long-term performance, highlighting the importance of robustness in the RLHF process.
@@ 6 Discussion
In this section, we provide insights into why the proposed reward-robust framework is effective and how it improves upon the previous standard pipeline. Given that the reward signal is inherently imperfect, we empirically demonstrate that over-scoring is more detrimental than under-scoring, which supports our choice of minimum return as a robustness measure. Additionally, we show that in the stochastic case where rewards are given randomly, selecting the minimum reward ensures that the trained model remains at least acceptable. We also do ablation study to compare our method with other integration strategies such as using mean reward in Section 6.3.
@@
We have demonstrated that the reward model is inherently imperfect (Section 4), meaning the reward used in the PPO training is either over-scored or under-scored. Through PPO training on minimized reward, maximized reward, and mean reward, we show that over-scoring is significantly more harmful than under-scoring in LLM RLHF. The results of this comparison are illustrated in Figure 4(a). In 12 of the total 16 benchmarks, the under-scoring setting outperforms the over-scoring setting (see Table 4).
@@
In our reward-robust RLHF framework, rewards tend to be under-scoring. To see this, we test BRME on UltraFeedBack Cui et al. (2024), a dataset with annotated scores. We approximately treat the annotation as the ground-truth reward, measuring the scoring gap of each reward head as well as the minimized reward head and the ground-truth reward. Note that the rewards are normalized with respect to its resource. The result is shown in Figure 5(b). It can be seen that for one single head, under-scoring and over-scoring are roughly equal. But in the reward-robust RLHF setting, minimizing rewards can make under-scoring the majority.
We give the following reasons for these observations: 1) In reward-robust RLHF, the optimization process tends to be more conservative, which could benefit long-term exploration. In essence, reward-robust RLHF might prioritize optimizing the lower bound of the return function. 2) In language tasks, exploring an incorrect direction may be more detrimental than rejecting a correct one. For a given prompt, there are often multiple valid responses. Even if the model misses one correct optimization path, it may still be able to explore alternative directions. However, with a more aggressive optimization strategy (such as over-scoring), the model might be more prone to pursuing a wrong direction, potentially resulting in reward hacking.
@@ 6.2 Stochastic Case Analysis
With a poorly-trained reward model, the worst-case scenario is that all rewards are given randomly. In this section, we demonstrate that even in such a stochastic case, the reward-robust RLHF framework still yields an acceptable model, and we provide an explanation for why this is the case. First, we have the following lemma,
@@
If we use reward robust RLHF training (setting $\lambda=0$ in Equation 7), as we are choosing the minimum reward, the range for reward will be narrowed (see Figure 5(b) and Figure 7(b)) and is closer to the constant reward situation. Since the starting point of our training is usually a well-trained SFT model, even if the optimization process degrades, the model can still maintain relatively good performance. In the contrast, if we use a random reward model, the optimization will be uncontrollable and the model will collapse in a fast speed.
@@
In the practice, the model optimization process is likely to oscillate, but it will eventually converge back to the inital-SFT model in a long run theoretically. The reason why it will oscillate is that the critic model is not ideal in the early stage, so the advantage $\widehat{A}_{t}$ is not 0 even all the rewards given are identical. However, as the KL divergence (-$\beta\log\frac{\pi_{\theta}(a|x)}{\pi_{0}(a|x)}$) is also considered in the real optimization objective (4), the actor policy will be eventually pulled back close to the initialized SFT policy.
@@ 6.3 Abalation Study
Our framework has two main different parts with the previous RLHF method. One is the Bayesian ensembled reward model. In previous works, the reward model directly output a scalar to be the final reward and the loss function is MLE(3). Another is the integration strategy. We use the trade-off version between the lower-bound reward and a nominal reward, while there are other strategies such as using the mean reward. Here we provide the ablation study results.
Comparison with reward model trained with MLE loss.
We trained another reward model using MLE loss on the same training data. The resulting single-head reward model is compared with the nominal head of the Bayesian ensemble reward model. To evaluate performance, we tested the accuracy on the preference dataset and compared their coverage. Additionally, we conducted PPO training using both reward models to compare their effects on the actors’ performance. The detailed results are shown in Appendix C.3.
@@
We conducted the same PPO experiment using an average of all reward signals and compared it with the min/max integration strategies. Additionally, we evaluated the trade-off strategy against the mean reward approach. The result is shown in Figure 9(c) and Figure 9(b). The performance of the mean strategy falls between the max and min strategies, but in the later stages of training (after 800 steps), the mean strategy also tends to lead to a decline in model performance. On most datasets where PPO has a significant effect, the reward-robust RLHF setting with a trade-off parameter of $\lambda=0.6$ outperforms the mean strategy.
@@ 7 Conclusions and Future works
In this paper, we proposed a reward-robust RLHF framework to address the problem of reward hacking in LLM alignment. We demonstrated that imperfection is an inherent characteristic of current reward model training, leading to the model exploring incorrect optimization directions. To mitigate this issue, we trained a Bayesian ensemble reward model with multi-head outputs to model the uncertainty set of the reward function. We showed that the head with the minimum variance effectively models the nominal reward function, representing the most confident scoring output. The newly proposed robustness-performance trade-off objective was proven effective, consistently outperforming baselines across most benchmarks. Furthermore, we demonstrated that under-scoring is preferable to over-scoring when dealing with imperfect reward models. Even in the worst-case scenario, where rewards are assigned randomly, the reward-robust RLHF framework still yields an acceptable model.
Since our method can be easily integrated into existing pipelines, there is potential to further improve performance by incorporating additional reward sources to better model the uncertainty set. Future work will explore the adoption of heterologous reward sources, including reward models trained on diverse datasets and direct scoring from closed-source LLM APIs such as GPT-4, as well as other markers mentioned in Yan et al. (2024a); Liu et al. (2023). The advantage of using heterologous models lies in their diverse base training datasets, which result in more varied reward scores, thereby improving the coverage of the uncertainty set. Preliminary exploration results on heterologous reward fusion are provided in Appendix F.
@@ Appendix A Annotator Disagreement in RLHF and RLAIF
For human-annotated data, we conducted an agreement test on 209 data points, each containing one prompt and two responses. The prompts were selected from an internal dataset by the PM team, and the responses were generated by Baichuan2-13B Yang et al. (2023). Data categories include general knowledge, logical reasoning, tables, mathematics, etc. We used two distinct annotator groups: one consisted of highly educated internal annotators who had undergone multiple rounds of specialized annotation training, referred to as the Expert Group. The other group was composed of external annotators hired from the general public, referred to as the External Group. For each data point, annotators were tasked with a Good-Same-Bad evaluation: 1) G: response 1 is better than response 2. 2) S: both responses are of the same quality. 3) B: response 1 is worse than response 2. We compared the G/S/B annotations for the same data between the two groups. Across all samples, the consistency rate was 70%, with a 4.5% rate of opposite judgments (one group scored G while the other scored B). When only considering G/B samples, the consistency rate increased to 77%, with an opposite rate of 6.5%.
@@
Appendix B Bayesian Reward Model Ensembles
In this section, we will provide the training detail, the theoretical explanation, and the empirical performance of the proposed Reward Model Ensembles (BRME).

### v3

@@ 5.1 BRME: Uncertainty Set Modeling
Figure 2: Diagram for BRME. Each head outputs the mean and the std of the corresponding reward distribution and reparametrization is emplyed to address the non-differentialbility.
In this section, we will show how we model the uncertainty set $\mathcal{R}^{\text{uncertain}}$ as well as the nominal reward function $\widehat{r}$. We propose Bayesian Reward Model Ensembles (BRME): We train a multi-head Bayesian reward model where the reward is modeled as a Gaussian distribution. Each head $i$ has two outputs: one representing the mean and the other representing the std. A sample from this distribution is then output as the reward.
The diagram for BRME is shown in Figure 2. To facilitate deployment and conserve computing resources, we do not train multiple RMs independently as an ensemble; instead, we use parameter sharing. All reward heads share a common base model, which serves as a feature extractor. During training, each data sample is randomly assigned to a head for training, with reparameterization employed to address the non-differentiability of the sampling process.
The training process is divided into two stages. In the first stage, we train a traditional one-head RM following the Maximum Likelihood Estimation (MLE) loss in (3). In the second stage, we leverage a Mean Squared Error (MSE) loss (8) to train the RM, which is first introduced in (Wu et al., 2024). The use of MSE loss to train the BRME ensures that: 1) the output’s std reflects the confidence of the model (see Appendix B.1), 2) the scoring range of each reward head is consistent. During prediction, each sample is evaluated by all heads, and the mean output by the head with the smallest std is selected as the nominal reward. We defer the thorough training pipeline, the theoretical analysis and detailed performance evaluation of BRME to Appendix B.
@@ 5.2.1 Experimental setup
We use LLaMa3-8B-Instruct (Dubey et al., 2024) as the initialization of the actor model. In BRME setting, we train a single-head RM and a 5-head BRME starting from LLaMa3-8B-Instruct (Dubey et al., 2024) as is described in Appendix B. The single-head RM is used as both the first stage model for BRME training and the initialization of the critic model. BRME is used solely as the reward signal source in PPO.
@@
For the training process, we use HH-RLHF (Bai et al., 2022a) and UltraFeedBack (Cui et al., 2024) to train the BRME. HH-RLHF, UltraFeedBack (Cui et al., 2024) , along with an internal prompt dataset collected by the PM team, is employed to implement the PPO algorithm. Details of the datasets are deferred to Table 6 in Appendix D. For performance evaluation, we select ARC (Clark et al., 2018), LAMBADA (Paperno et al., 2016), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), WinoGrande (Sakaguchi et al., 2019), TQA (Lin et al., 2022), MMLU (Hendrycks et al., 2020), GSM8K (Cobbe et al., 2021), FDA (Arora et al., 2023), EQ-Bench (Paech, 2023), Arithmetic (Brown et al., 2020), and ANLI (Nie et al., 2020) as our benchmarks. The evaluation dimensions include robustness, general knowledge, numerical computation, emotion reading, information extraction, reasoning, context understanding, and commonsense. Detailed descriptions of the datasets are shown in Table 7 in Appendix D.
@@
The RM training and PPO experiments are conducted on 24 Nvidia H800-SXM-80GB GPUs in 3 nodes using DeepSpeed library, ZeRO stage 2 (Rasley et al., 2020), and HuggingFace Accelerate (Gugger et al., 2022). In PPO process, the actor model and the critic model occupy 10 gpus respectively. The reference model and BRME occupy 2 gpus respectively. We use AdamW optimizer (Loshchilov et al., 2017). The experience batch size in PPO is set to be 128.
@@ 5.2.2 Main Results
We incrementally increase the performance trade-off hyperparameter $\lambda$ from 0 to 1 in intervals of 0.2 and repeated PPO training under each setting for 800 steps. It is important to note that when $\lambda=1$, the algorithm degrades to standard RLHF with a single nominal RM. The performance evaluation results at 200 and 800 steps are shown in Figure 3.
@@
In the short run (200 steps), although there are a few exceptions—such as ANLI-r1 and LAMBADA, where $\lambda=0$ outperforms the other settings—in most cases, the trade-off versions with $\lambda=0.4$ and $\lambda=0.6$ show better performance. This suggests that, even early in training, incorporating a balance between performance and robustness offers notable advantages. The ability of the reward-robust RLHF framework to temper the optimization process appears to result in more stable performance gains compared to standard RLHF. Compared with standard RLHF, the average accuracy of reward-robust RLHF increases by 0.99% and 1.40% respectively when $\lambda=0.4$ and $0.6$. On certain datasets, such as arithmetic and ANLI-r3, the improvement exceeds 3%.
@@
Over the long run (800 steps), the advantages of incorporating robustness measurement become even more pronounced. Nearly all experimental groups with $\lambda\neq 1$ exhibit improved performance by the end of 800 steps, confirming the long-term benefits of reward-robust RLHF. In contrast, standard RLHF with $\lambda=1$ not only fails to capitalize on the full optimization process but, in some cases, even results in negative performance growth. For example, in tasks like ANLI-r1 and GSM8K, we observe a decrease in model capability during the additional 600-step optimization process under standard RLHF. This highlights a key limitation of the purely performance-driven RLHF approach: its vulnerability to the risk of overfitting or misguidance from imperfect reward signals. Compared with standard RLHF, the accuracy of reward-robust RLHF increases by 2.42% and 2.03% respectively when $\lambda=0.4$ and $0.6$.
@@
On tasks where the RM performs notably poorly, such as MMLU, we notice that PPO training under standard RLHF can degrade model performance, evidenced by a drop in accuracy, which is reflected in the bar graphs pointing downward. This degradation likely results from the model being optimized based on unreliable or misleading reward signals, which causes the policy to drift away from optimal behavior. In contrast, the reward-robust RLHF approach, which balances performance with robustness ($\lambda=0.4$ or $\lambda=0.6$), mitigates this decline. By narrowing the optimization focus and stabilizing the reward signal, the model is better able to resist overfitting to faulty rewards, ultimately preserving and improving performance. The results suggest that in settings where training data is sparse and the RM is significantly imperfect, adopting a reward-robust approach can effectively stabilize training and prevent further performance degradation over time.
@@ 6 Discussion
In this section, we provide insights into why the proposed reward-robust RLHF framework is effective and how it improves upon the previous standard pipeline. Given that the reward signal is inherently imperfect, we empirically demonstrate that over-scoring is more detrimental than under-scoring, which supports our choice of minimum return as a robustness measure. Additionally, we show that in the stochastic case where rewards are given randomly, selecting the minimum reward ensures that the trained model in PPO remains at least acceptable. In Section 6.3, we also do ablation study to compare our method with traditional RM with MLE loss as well as other integration strategies such as using mean reward.
@@
We have demonstrated that RMs are inherently imperfect (Section 4), meaning the rewards used in the PPO training are either over-scored or under-scored. Through PPO training on minimum reward and maximum reward, we show that over-scoring is significantly more harmful than under-scoring in PPO training. The results of this comparison are illustrated in Figure 4(a). In 12 of the total 16 benchmarks, under-scoring setting outperforms the over-scoring setting (see Table 4, Appendix C.2).
@@
In our reward-robust RLHF framework, rewards tend to be under-scoring. To see this, we test BRME on UltraFeedBack (Cui et al., 2024), diverse preference dataset with annotated rewards. We approximately treat the annotation as the ground-truth reward, measuring the scoring gap of each reward head as well as the minimum reward head and the ground-truth reward. Note that the rewards are normalized with respect to its resource. The result is shown in Figure 5(b). For a single head, the number of under-scoring and over-scoring cases is roughly equal. However, in the reward-robust RLHF setting, selecting the minimum rewards leads to under-scoring becoming more frequent.
We propose the following hypotheses to explain these observations: 1) In reward-robust RLHF, the optimization process tends to be more conservative, which could benefit long-term exploration. In essence, reward-robust RLHF might prioritize optimizing the lower bound of the return function. 2) In language tasks, exploring an incorrect direction may be more detrimental than rejecting a correct one. For a given prompt, there are often multiple valid responses. Even if the model misses one correct optimization path, it may still be able to explore alternative directions. However, with a more aggressive optimization strategy (such as over-scoring), the model might be more prone to pursuing a wrong direction, potentially resulting in reward hacking.
@@ 6.2 Stochastic Case Analysis
If the RM is poorly trained or handling OOD data, the extreme scenario is that rewards are given randomly. In this section, we demonstrate that even in such a stochastic case, the reward-robust RLHF framework still yields an acceptable model, and we provide an explanation for why this is the case. First, we have the following lemma,
@@
If we use reward robust RLHF training (setting $\lambda=0$ in Eq. 7), as we are choosing the minimum reward, the range for reward will be narrowed (see Figure 5(b) and Figure 8(b)) and is closer to the constant reward situation. Since the starting point of our training is usually a well-trained SFT model, even if the optimization process degrades, the model can still maintain relatively good performance. In the contrast, if we use a random reward model, the optimization will be uncontrollable and the model will collapse in a fast speed.
@@
In the practice, the model optimization process is likely to oscillate, but it will eventually converge back to the inital-SFT model in a long run theoretically. The reason why it will oscillate is that the critic model is not ideal in the early stage, so the advantage $\widehat{A}_{t}$ is not 0 even all the rewards given are identical. However, as the KL divergence (-$\beta\log\frac{\pi_{\theta}(a|x)}{\pi_{0}(a|x)}$) is also considered in the optimization objective (4), the actor policy will be eventually pulled back close to the reference policy.
@@ 6.3 Abalation Study
Our framework has two main different parts with the previous RLHF method. One is the BRME setting trained with MSE loss. In most previous works, the RM directly output a scalar to be the final reward and the loss function is MLE(3). Another is the integration strategy. We use the trade-off version between the lower-bound reward and a nominal reward, while there are other strategies such as using the mean reward Eisenstein et al. (2023). Here we provide the ablation study results.
Comparison with RM trained with MLE loss.
We trained another RM using MLE loss on the same training data as mentioned in Appendix B. The resulting single-head RM was then compared with the nominal head of the BRME. To assess performance, we evaluated the accuracy on a series of preference datasets, including General QA, Writing, Comprehension, and Math (Table 5). One key advantage of BRME is its ability to effectively model the diversity within the uncertainty set, which directly influences optimization Mannor et al. (2016). To demonstrate this, we also compared the reward signal distribution coverage (Figure 7). Additionally, we conducted PPO training using both RMs to compare their effects on the actors’ performance. The results consistently support BRME’s superiority, with detailed findings provided in Appendix C.3.
@@
We conducted the same PPO experiment using an average of all BRME heads and compared it with the min/max integration strategies. Additionally, we evaluated the trade-off strategy against the mean reward approach. The result is shown in Figure 9(c) and Figure 9(b), Appendix C.4. The performance of the mean strategy falls between the max and min strategies, but in the later stages of training (after 800 steps), the mean strategy also tends to lead to a decline in model performance. On most datasets where PPO has a significant effect, the reward-robust RLHF setting with a trade-off parameter of $\lambda=0.6$ outperforms the mean strategy.
@@ 7 Conclusions and Future works
In this paper, we proposed a reward-robust RLHF framework to address the problem of reward hacking in LLM alignment. We demonstrated that imperfection is an inherent characteristic of current reward model training, leading to the model exploring incorrect optimization directions. To mitigate this issue, we trained BRME with multi-head outputs to model the uncertainty set of the reward function. We showed that the head with the minimum std effectively models the nominal reward function, representing the most confident scoring output. The newly proposed robustness-performance trade-off objective was proven effective, consistently outperforming baselines across most benchmarks. Furthermore, we demonstrated that under-scoring is preferable to over-scoring when dealing with imperfect RMs. Even in the stochastic-case scenario, where rewards are assigned randomly, the reward-robust RLHF framework still yields an acceptable model. Finally, we acknowledge the framework’s limitations, which are discussed in detail in Appendix G.
Since our method can be easily integrated into existing pipelines, there is potential to further improve performance by incorporating additional reward sources to better model the uncertainty set. Future work will explore the adoption of heterologous reward sources, including RMs trained on diverse datasets and direct scoring from closed-source LLM APIs such as GPT-4, as well as other markers mentioned in (Yan et al., 2024a; Liu et al., 2023). The advantage of using heterologous models lies in their diverse base training datasets, which result in more varied reward scores, thereby improving the coverage of the uncertainty set. Preliminary exploration results on heterologous reward fusion are provided in Appendix F.
@@ Appendix A Annotator Disagreement in RLHF and RLAIF
For human-annotated data, we conducted an agreement test on 209 data points, each containing one prompt and two responses. The prompts were selected from an internal dataset by the PM team, and the responses were generated by Baichuan2-13B (Yang et al., 2023). Data categories include general knowledge, logical reasoning, tables, mathematics, etc. We used two distinct annotator groups: one consisted of highly educated internal annotators who had undergone multiple rounds of specialized annotation training, referred to as the Expert Group. The other group was composed of external annotators hired from the general public, referred to as the External Group. For each data point, annotators were tasked with a Good-Same-Bad evaluation: 1) G: response 1 is better than response 2. 2) S: both responses are of the same quality. 3) B: response 1 is worse than response 2. We compared the G/S/B annotations for the same data between the two groups. Across all samples, the consistency rate was 70%, with a 4.5% rate of opposite judgments (one group scored G while the other scored B). When only considering G/B samples, the consistency rate increased to 77%, with an opposite rate of 6.5%.
@@
Appendix B Bayesian Reward Model Ensembles (BRME)
In this section, we will provide the training detail, the theoretical explanation, and the empirical performance of the proposed Bayesian Reward Model Ensembles (BRME).